OpenStudy (anonymous):

how do I convert 55 mph to ft/s

OpenStudy (lgbasallote):

i can teach you two ways...the google way (they are magnificent in conversion) or the self-solving way which do you prefer?

OpenStudy (anonymous):

the self-solvingway

OpenStudy (lgbasallote):

okay... \[55 mph \implies \frac{55 \; \text{miles}}{1 \; \text{hour}}\] agree?

OpenStudy (lgbasallote):

that's what mph means right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now... do you agree that in 1 hour there are 3,600 seconds?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now i write.. \[\frac{55 \; \text{miles}}{1 \; \text{hour}} \times \frac{1 \; \text{hour}}{3,600 \; \text{seconds}}\] still following?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

now... in 1 mile, there are 5,280 feet so... \[\frac{55 \; \text{miles}}{1 \; \text{hour}} \times \frac{1 \; \text{hour}}{3.600 \; \text{seconds}} \times \frac{5,280 \; \text{feet}}{1 \; \text{mile}}\] do you still get it?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now, we cancel the units \[\frac{55 \; \cancel{\text{miles}}}{1 \; \cancel{\text{hour}}} \times \frac{1 \; \cancel{\text{hour}}}{3.600 \; \text{seconds}} \times \frac{5,280 \; \text{feet}}{1 \; \cancel{\text{mile}}}\] so now, just simplify that with a calculator and you get your answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

60mph = 88 ft/sec 55/60 of 88 is 80.7 ft/sec rounded to the nearest 1/10